Operations Report – November 25th

Crew 216 Operations Report 25-Nov-19
SOL: 1
Name of person filing report: Mike Lawson
Non-nominal systems:
Notes on non-nominal systems: N/A

Generator: Run
Hours run: 12
From what time last night: 1909
To what time this morning: 0710
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night: 72% as of 1649
Diesel Reading – 43
Station Propane Reading – 61
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 47
Water Meter (units): 1461353
Water (static tank) (gal): 436
Static to Loft Pump used – Yes
Water in Green Hab (gal): 279
Water in Science Dome (gal): 0
Toilet tank emptied: Yes

Deimos rover used: No
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: Yes
Hours: 114
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 65
Currently charging: Yes

Opportunity rover used: Yes
Hours: 69.7
Beginning charge: 100
Ending charge: 50
Currently charging: Yes

Curiosity rover used: Yes
Hours: 119
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 42
Currently charging: Yes

Notes on rovers: 100

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs:

HabCar used and why, where? N/A
CrewCar used and why, where?

General notes and comments: Crew morale is high.
Summary of internet: Crew usage triage meeting tonight.
Summary of suits and radios: Conducted a ranging test between EVA and HAB, comparing Midland H/H and TERA TR-590 with 5/8th wave antenna. Superior performance from the TERA; will use for future EVAs as HAB NCS on Channel 22.
Summary of Hab operations: Repaired exposed wiring connection in HAB living area.
Summary of GreenHab operations: Routine watering.
Summary of Science Dome operations: Gypsum samples deposited for further processing. Agate and pretrified wood samples UV tested. Fluoresed yellow/red and magenta respetively under low band UV.
Summary of RAM operations: Require replacement light bulb, standard screw in base. Airlock ready room bulb broken, have cannibalized bulb from spare bedroom. 0 balance on hand. Fabricated expedient 4 x 4 junction box cover, exposed wiring connections, HAB living area. 0 balance o/h.
Summary of any Observatory issues: No solar obseration because of cloud cover.
Summary of health and safety issues: Placed CO alarm is spare stateroom. Some crew complaining of evening headaches. Likely just symptomatic of mild dehydration, test is simply to eliminate a possible cause.

Questions, concerns and requests to Mission Support: Light bulb resupply. Suggest two, one for spare stateroom and one for maintenance float. Need power cable (outlet to transformer) for astronomy laptop.

Copyright © The Mars Society. All rights reserved. | Main Site